OpenStudy (anonymous):

An investment is worth $3518 in 1995. By 2000 it has grown to $5553. Let y be the value of the investment in the year x, where x = 0 represents 1995. Write a linear equation that models the value of the investment in the year x. A. y = -407x + 7588 B. y = 1/407 x + 3518 C. y = -407x + 3518 D. y = 407x + 3518

OpenStudy (jdoe0001):

so, in year 1995, it was $3518, and in 2000, 5 years later, $5553, how much has it grown?

OpenStudy (campbell_st):

so the y intercept of the line is when x = 0... $3518 and after 5 years you have $5553 so to find the slope its m = (5553 - 3518)/(5 - 0) the sline will be y = mx + 3518

OpenStudy (anonymous):

so is it D @campbell_st

OpenStudy (campbell_st):

looks like it

OpenStudy (anonymous):

ok thank you so much for your help

OpenStudy (anonymous):

you too @jdoe0001

OpenStudy (jdoe0001):

yw